# Revitalising the Single Batch Environment: A ’Quest’ to Achieve Fairness and Efficiency

### v2

@@ 3.3 Analysis of the Batch: Best and Worst Cases
In our setting, SPT and SRPT is equivalent in terms of working and performance
@@
SRTF switches the context iff at any instance it discovers some lower burst(s) compared to the process currently under execution. In our setting, we have all the processes in the batch available from the very beginning and there is no addition of other jobs at any point in time. SRTF starts with the least burst from the batch and after each unit of time, the remaining time of the least burst is trivially the least as there is guaranteed progress which is why there is no preemption as there cannot be any other job at any instance which would have lesser burst than the currently executing process for all jobs in the batch. This is exactly the way SPT works. ∎
@@
remark: For the sake of consistency, we have used SRTF in lieu of SPT throughout manuscript.
Maximum Average Response Time for the Batch
@@
This is considerably the worst case for response time. The batch can have the maximum waiting time when all the waiting processes have to wait the most. As exactly ’n-1’ jobs from a batch with ’n’ jobs will wait whenever 1 job is under execution, if the currently executing job itself is the highest burst amongst the processes at each iteration and there is no preemption until the entire completion of the executing job, the rest of the jobs wait for the longest possible time. We can simulate to achieve this scenario using LPT. As it is known to maximise the average completion time [3], so it trivially maximises responsetime in a single server setting, just like ours.
Maximum Average Waiting Time for the Batch
@@
This is the worst case for waiting time. LPT, as highlighted in [3], stands as the inverse of the SPT policy, and maximises completion time. However, since the conventional approach prioritizes minimization, especially in our comparison with SRTF/SPT algorithms, we opt to evaluate the preemptive variant, LRTF. This choice aims to assess its potential for enhanced response times due to its preemptive nature.
@@
remark: As our main goal is to revitalise the single batch environment, we’ve given the results for the very extreme cases above but in real life, to the best of our knowledge, we’ve not found out any single algorithm which simultaneously achieves best of 2 world (such as best average waiting and response time or vice-versa). Any algorithm including FairBatch, designed for this setting, would always fall within the extreme ranges demonstrated. As the concentration for this research is on considering both waiting and response time together across distribution. The trivial best case for both waiting and response time achieved when simply the batch has 1 job using proposition 3.1. However algorithmic behavior changes and shows different alignment with regards to the underlying distribution. Though the best case is trivial but the worst case where both waiting and response time is maximised is not achievable as the algorithm is simply not LRT. As a result, the worst case comes with a certain trade-off between response and waiting time which is always distribution dependent, although individually average waiting and average response times would be lie within the ranges derived for the batch. This is why, we’ve studied and investigated exhaustively how our algorithms behave under out robust experimental setup. Now, there can be so many types of distributions, for our research, we have considered those which are commonly observed in real-life scenarios the most and can simulate the problems that we had mentioned earlier such as convoy and starvation effects etc.
3.4 Experimental Setup
FCFS is the most commonly used algorithm for batches but we’ve also taken preemptive and advanced algorithms like SRTF, LRTF, RR along with FCFS for comparison. In our comprehensive study on batch processing algorithms, we aimed to provide a robust evaluation by considering diverse bursttime distributions and a large number of test cases. We carefully selected six principle probability distributions - Normal, Exponential, Geometric, Negative Binomial, Poisson, and Uniform - to generate bursttime data and experimented on 3 multi-model distributions later. The selection of six diverse probability distributions allowed us to explore various bursttime patterns and behaviours. Each distribution served a unique purpose in our study. The selection of the algorithms is mainly based on the cases where the distribution of the bursts is expected to cause convey effect and/or starvation. This is due to the reason for inspecting the relative behaviour of our algorithm in these situations along with other distributions. The Normal distribution enabled us to assess algorithms’ ability to handle data clustering around a mean value, while the Exponential distribution shed light on performance in scenarios with exponential decay-like patterns. The Geometric distribution offered insights into algorithms’ response to decreasing probabilities of longer bursttimes. Simulating long-tailed variations, the Negative Binomial distribution, being a general case of the geometric distribution, allowed us to scrutinize algorithm behaviour in such scenarios. By employing the Poisson distribution, we observe algorithm behaviour under specific occurrence patterns of processes. Lastly, the Uniform distribution challenged algorithms with bursttimes exhibiting a uniform and evenly distributed pattern.
Figure 2: Results across Distributions
Outliner of the testcases
For each distribution, we generated 100 test cases, with each test case consisting of 100 integers bursttimes. The bursttimes were generated following the properties of the respective distribution, with values ranging from 1 to 500. For each list, we have scaled the data in such a way that after scaling, the sum of bursttimes in each and every list are equal or very close to equal. For illustration here, we’ve taken the sum of the bursttime in each list to be approximately equal to 25,000. By scaling the data, we eliminate the influence of magnitude disparities, allowing us to focus solely on the algorithmic performance with respect to bursttime distribution patterns. For a fine-grained comparison, we’ve used box plots for each metric to analysis across various distributions for each algorithm. We’ve referred to our algorithm as ”FairBatch” here and have taken the ceil of mean of the bursttimes as the timeQuantum in RR as it remains almost equal for all the lists under examination. In the diagrams ’Nor Distribution’ stands for Normal Distribution, ’Expo Distribution’ stands for Exponential Distribution, ’Geo Distribution’ stands for Geometric Distribution, ’N.Bio Distribution’ stands for Negative Binomial Distribution, ’Psn Distribution’ stands for Poisson Distribution, ’Ufm Distribution’ stands for Uniform Distribution, 1 stands for FCFS, 2 stands for FairBatch, 3 stands for SRTF, 4 stands for LRTF, 5 stands for RR.
Figure 3: Results across Distributions
Figure 4: comparison across multimodal distribution
3.5 Discussion
First and foremost, in designing the algorithm, our primary focus was on achieving a ’holistic’ selection of processes at each round of execution comparing the fairnessRatio we engineered. Figure 2 consists of exponential, geometric, and negative binomial distributions, figure 3 consists of Poisson, normal, and uniform distributions respectively. This careful selection is the key factor that allows the algorithm to outperform others in terms of average response time in most of the cases including the ’responsive’ algorithm round-robin. The algorithm consistently prioritizes ’holistically deserving’ processes at each cycle, leading to a substantial performance advantage over other algorithms, except for one interesting case.
Investigating the performance in the case of Normal Distribution
In the case of normal distribution in LRTF, we can observe that the average response time is much better than FairBatch(Figure 3). In LRTF scheduling policy, when the processes are clustered around the mean in a normal distribution, there tends to be a substantial cluster of potentially longer processes. LRTF prioritizes these long processes rapidly, giving them higher preference over shorter ones during scheduling. As a result, the long processes get executed sooner, reducing the overall response time. However, this rapid prioritisation of long processes in LRTF comes at a cost – other processes starve for execution. We can see this in the waiting time plots. In contrast, our FairBatch algorithm effectively eliminates starvation using its fairnessRatio and sorting mechanism with the dynamic time slice, as evident in the charts.
However, there is a case where FCFS excels and FairBatch falls short. In the case of normal distribution, FCFS performed better in terms of average turnaround time in Figure 3. Being a non-preemptive algorithm FCFS doesn’t switch the context while executing a cluster of processes around the mean of the bursttime. Causing less responsiveness due to this convoy effect. FairBatch efficiently mitigates the convoy effect with its optimal trade-off between waiting and response time. Notably, this trade-off substantially outperforms ’responsive’ algorithms like round robin which is evident from the box plots.
Observing the trade-off Arose in the case of Poisson Distribution
Poisson distribution is mainly used to show how many times an event is likely to occur over a specified period. It’s clear from the plots that the algorithm, unlike others, is not primarily focused on minimising one parameter rather, it’s design elegantly establishes a trade-off between waiting and response time. FairBatch excels in all other cases including the uniform distribution, which closely resembles the real-world scenarios, across all algorithms, outperforming its counterparts in terms of all three metrics, except when compared to the benchmark in waiting and turnaround times. This demonstrates that the FairBatch algorithm exhibits strong adaptability and resilience in handling the randomness present in different data distributions.
Further Investigating the Trade-off
The strength and applicability of an algorithm stem from its ability to strike a balance in computationally expensive scenarios. Greedy scheduling policies, such as SRTF, may minimize waiting time but at the cost of reduced responsiveness followed by starvation. Conversely, algorithms like LRTF prioritize longer tasks, which may overlook other important factors.
To further demonstrate the significance of this trade-off, we’ve considered three multi-modal distributions in Figure 4 by concatenating distinct normal data sets. In the first two cases of bimodal and trimodal distributions, we’ve arranged non-decreasing clusters in the dataset. The charts clearly show that the waitingtime and turnaround time of FCFS and SRTF are comparable in both cases. As the number of local maxima in a non-decreasing multi-modal dataset increases, the difference between these two algorithms diminishes due to randomisation. This can be a case of switching algorithms by the scheduler ’S’ as discussed in the proposal of the work section. Although SRTF, the benchmark, still minimized the waiting and turnaround time, it cannot maintain a responsive flow for the process due to its high response time causing starvation. FairBatch, in all these cases where there exist different types of sub-populations or clusters of processes, prioritizes the trade-off discussed earlier. From the charts, it is evident how FairBatch efficiently schedules the processes, removing starvation while maintaining a competitive range for turnaround and waitingtime. In the multi-modal distribution, we have taken a shuffled permutation of different sub-populations, and here also FairBatch successfully mitigates the major issues across various algorithms, establishing a solid trade-off, proving its universality across numerous alignments of data.
Stability of the performance: another advantage of FairBatch
FairBatch offers a distinct advantage over algorithms like Round Robin and FCFS due to its stability and predictability. While the sequence of execution and overall results of these algorithms heavily depend on the sequence in which the job arrived, FairBatch eliminates this variability by arranging and executing jobs based on their inverse proportion of bursttime and the designated ratio metric from the very first iteration. This characteristic makes FairBatch a more reliable and consistent scheduling algorithm, ensuring stability and predictability in its performance.
Concluding thoughts:
FairBatch demonstrates its superior performance across various data distributions, outperforming other algorithms in most cases. Notably, FairBatch competes competitively with the benchmark algorithm, SRTF, in terms of average turnaround as well as waiting time and surpasses it in responsiveness.
After a comprehensive comparison of all algorithms across various distributions as demonstrated in Figure 2, Figure 3 and Figure 4, it is evident that the FairBatch performs exceptionally competitively with respect to the benchmarks for all distributions showcasing its superiority through the delicate balance it strikes between drastically reducing waitingtime, ensuring responsiveness, and eliminating starvation. With stability, predictability, and unmatched efficiency, our algorithm bridges the gap between suitable responsiveness and required efficiency through a fair and holistic approach, as demonstrated extensively.
4 Optimisation of the Algorithm
@@
This ’Select’ algorithm outputs a set of appropriate proportions of job(s) based on fairnessRatios and timeQuantum. The ’Runner’ algorithm takes the set of processes ’s’ as a input from ’Select’. It sorts the processes in ’s’ based on their fairnessRatio. Run the processes sequentially and update their attributes. After the cycle ends, the attributes of the process not were under execution is updated and the scheduler runs the subsequent cycles till there is process left in the batch.
@@
The ’Select’ algorithm starts by taking as many processes with values greater than the median as possible, staying within the weight constraint. If there is any leftover capacity, it takes processes with values equal to the median. If further capacity remains, the algorithm recursively considers combinations of processes to maximize the total value while respecting the timeQuantum limit. The median calculation, utilized in both FairBatch and the ’Select’ procedure, can be computed in linear time. In ’Select’, each recursive call requires linear time, excluding the time spent on potential recursive calls it may make. Since there is only one recursive call, it pertains to a problem size at most half of the original. As a result, the running time can be expressed by the following recurrence relation: T(n) <=<=< = T(n/2)+ $\theta$(n) therefore, T(n) = O(n), using master’s theorem.
@@
If the ’Select’ algorithm returns ’k’ jobs to be scheduled, ’Runner’ will take additional k.log k asymptotic time complete the scheduling due to sorting. So overall the scheduler takes: O(n+ k.log k) per cycle where k <=<=< = n.

### v4

@@ 3.3 Further Analysis of the Parameters:
    In our setting, SJF and SRTF are equivalent in terms of working and performance
@@
    SRTF switches the context iff at any instance it discovers some lower burst(s) compared to the process currently under execution. In our setting, we have all the processes in the batch available from the very beginning and there is no addition of other jobs at any time. SRTF starts with the least burst from the batch and after each unit of time, the remaining time of the least burst is trivially the least (as there is guaranteed progress using proposition 3.1). This is why there is no preemption as there cannot be any other job at any instance which would have a lesser burst than the currently executing process for all jobs in the batch. This is exactly the way SJF works. ∎
@@
    remark: For the sake of consistency, we have used SRTF and SJF interchangeably throughout the manuscript.
  • Maximum Average Response Time for the Batch:
@@
    In our setting, the batch will experience the maximum response time when all the waiting processes have to wait the most. Exactly ‘n-1’ jobs from a batch with ‘n’ jobs will wait whenever exactly 1 job is under execution. If every time the currently executing job itself is of the highest burst amongst the processes and there is no preemption until the entire completion of the executing job, the rest of the jobs wait for the longest possible time. Longest Process First(LPT) chooses the longest job to execute without preemption. As a result, in our environment for a single batch with a single processing unit, it trivially maximises the responsetime. It also maximises the average completion time [30].
  • Maximum Average Waiting Time for the Batch:
@@
    LPT is the inverse of the SJF policy and maximises mean completion time [30]. However, since the conventional approach prioritizes minimization of completion time, especially in our comparison with SRTF/SJF algorithms, we opt to evaluate the preemptive variant, LRTF. This choice aims to assess its potential for enhanced response times due to its aggressive preemptive nature.
@@
remark: As our main goal is to revitalise the single batch environment, we’ve given the results for the extreme cases above but in realistic scenarios, to the best of our knowledge, there is no single algorithm that simultaneously achieves the best of 2 world (lowest possible average waiting and response time). Any algorithm including FairBatch, designed for this setting, would always fall within the extreme ranges demonstrated. As the concentration for this research is on considering both waiting, turnaround and response time together across the distribution, the trivial case for lowest both waiting and response time is achieved when simply the batch has 1 job using proposition 3.1. However algorithmic behaviour changes and shows different alignments with regards to the underlying distribution of burst time of the jobs.
This is why, we’ve studied and investigated exhaustively how our algorithms behave under our robust experimental setup in Section 4. Now, there can be many types of distributions, for our research, we have considered those that are commonly observed in real-world scenarios the most and can simulate the problems that we had mentioned earlier such as convoy and starvation effects etc.
3.4 Importance of a Suitable Time Quantum
In a queue of processes, even when executing just one process for the minimal feasible time, (say “t”) for the sake of achieving utmost fairness, we must re-calculate the fairness ratio of every process within the batch after each “t” unit of time. Now “t” mathematically can be arbitrarily small or large. If “t” is exceedingly small, process progress becomes exceedingly marginal with excessively unnecessary context switches, while an arbitrarily large “t” renders the algorithm’s behaviour akin to First-Come-First-Serve (FCFS) for the majority of the time – both of which contradicts the philosophy of FairBatch. Moreover in realistic scenarios, the distribution on job’s bursttimes may follow several distribution or might be completely random in nature. For example, in a positively skewed distribution, the mean is greater than the median and visa versa for negatively skewed counterpart. Regardless of distributions or random distributions, mean and median are always prominent factors to get an overview of the central tendency. This is why we have diligently explored diverse statistical formulations across current literatures [32, 33, 34, 35, 36, 37], and have finalised, inspired by the empirical work presented in [38].
3.5 Optimisation of the Algorithm
@@
This ‘Select’ algorithm outputs a set of appropriate proportions of job(s) based on fairnessRatios and timeQuantum. The ‘Runner’ algorithm takes the set of processes ‘s’ as a input from ‘Select’. It sorts the processes in ‘s’ based on their fairnessRatio. Run the processes sequentially and update their attributes. After the cycle ends, the attributes of the process not were under execution is updated and the scheduler runs the subsequent cycles till there is process left in the batch.
@@
The ‘Select’ algorithm starts by taking as many processes with values greater than the median as possible, staying within the weight constraint. If there is any leftover capacity, it takes processes with values equal to the median. If further capacity remains, the algorithm recursively considers combinations of processes to maximize the total value while respecting the timeQuantum limit. The median calculation, utilized in both FairBatch and the ‘Select’ procedure, can be computed in linear time. In ‘Select’, each recursive call requires linear time, excluding the time spent on potential recursive calls it may make. Since there is only one recursive call, it pertains to a problem size at most half of the original. As a result, the running time can be expressed by the following recurrence relation: T(n) <=<=< = T(n/2)+ $\theta$(n) therefore, T(n) = O(n), using master’s theorem.
@@
If the ‘Select’ algorithm returns ‘k’ jobs to be scheduled, ‘Runner’ will take additional k.log k asymptotic time complete the scheduling due to sorting. So overall the scheduler takes: O(n+ k.log k) per cycle where k <=<=< = n.
4 Experimental Setup
FCFS is the most commonly used algorithm for batch jobs but we’ve also taken preemptive and advanced algorithms like SRTF, LRTF, RR, and CFS along with FCFS for comparison. Along with theoretically well grounded policies such as SRTF, LRTF, RR and FCFS, we have considered adopting the implementation CFS algorithm ( [^1] https://elixir.bootlin.com/linux/v5.19.9/source/kernel/sched/fair.c#L7429) for a uni processing system as per our setting. We consider the default ‘nice’ value of individual jobs in the batch as 0222https://man7.org/linux/man-pages/man7/sched.7.html. By comparing with these traditional and contemporary scheduling approaches enriches the evaluation, blending theoretical robustness with practical relevance, and ensuring a comprehensive assessment.
Refer to captionRefer to captionRefer to caption
Figure 2: Exponential, Geometric & Negative Binomial clusters
Refer to captionRefer to captionRefer to caption
Figure 3: Normal, Poisson, & Uniform clusters
Refer to captionRefer to captionRefer to caption
Figure 4: Pareto, Gamma, & Standard Cauchy (Lorentz) clusters
Refer to captionRefer to captionRefer to caption
Figure 5: Bimodal, Trimodal, & Tultimodal clusters
4.1 Description of the Test Cases
Unlike recent works [32, 33, 34, 35, 36, 37] and others where primarily evaluations of algorithms are based on mere set of unsystematic, tiny examples, we aim to provide a robust evaluation by considering diverse set of clusters of jobs and a large number of test cases. To the best of our knowledge, there is no publicly available, sufficiently large, domain-agnostic, benchmark dataset of bursttimes that would be suitable for comparing under our setting and methodologies. It was a significant motivation for us to first create a diverse and sufficiently large dataset of bursttimes. We have made it publicly available with the aim of enabling the research community to utilize it in their own work. Eventually, we evaluated all algorithms on this dataset to assess their performance. For this evaluation, first of all we have carefully selected several probability distributions to generate the test cases: we have considered Normal, Exponential, Geometric, Negative Binomial, Poisson, Uniform, Pareto, Gamma, and Standard Cauchy (Lorentz) distribution. The Normal distribution enabled us to assess algorithms’ ability to handle data clustering around a mean value, while the Exponential distribution shed light on performance in scenarios with exponential decay-like patterns. The Geometric distribution offered insights into algorithms’ response to decreasing probabilities of longer burst times. Simulating long-tailed variations, the Negative Binomial distribution, being a general case of the geometric distribution, allowed us to scrutinize algorithm behavior in such scenarios. By employing the Poisson distribution, we observe algorithm behavior under specific occurrence patterns of processes. The Uniform distribution challenged algorithms with burst times exhibiting a uniform and evenly distributed pattern. The Pareto distribution, on the other hand, is characterized by a heavy tail, making it suitable for modeling phenomena where a small number of events have a disproportionately large impact. The Gamma distribution, with its shape and rate parameters, offers versatility in modeling various phenomena, including wait times and service times. The Standard Cauchy distribution, also known as the Lorentz distribution, represents a distribution with heavy tails and no defined mean or variance, providing insights into scenarios with extreme outliers. Together, these distributions provide a comprehensive overview of diverse bursttime patterns [39], offering a holistic assessment of algorithm performance across various real-world scenarios, which has not been covered in previous studies. While the distributions we’ve covered cover many scenarios, sometimes load on the CPUs can be even more varied and complex. To ensure our evaluations are robust, we’ve also considered bimodal, trimodal, and multimodal distributions. In a continues frame of reference (the CPU is processing arbitrarily large number of jobs),these distributions exhibit more irregular and diverse bursttime patterns. In the first phase of analysis we study the first nine distributions and in the second phase we investigate into these three special distributions.
For each distribution, we first produce 100 test cases, with each test case consisting of 100 integers bursttimes. After this initial phase, we fine tune the entire dataset. While distributions are primarily generated [^3] https://numpy.org/doc/1.16/reference/routines.random.html by randomly sampling values, these samples may not align with the realistic characteristics of job burst times, where each job’s burst time must fall within a feasible range. Simply relying on theoretical distributions could lead to unrealistic or unfeasible burst times, skewing the evaluation of scheduling algorithms. Our fine tuning are 2 phase:
  1. 1.
    Adhering to Practical Constraints: While we initially generated instances from several distributions to emulate diverse loads on the CPU, we recognize that CPUs do not process loads exactly the way they arrive in the system (in batch form here) following a precise distribution. CPUs cannot process arbitrarily small portions of a job; they allocate a minimum execution time for all jobs by default. Similarly, CPUs cannot allocate arbitrarily large processes. If a process exceeds the CPU’s allocation, the remainder of the process is stored in secondary memory. Using several page ranking algorithms [40], the CPU executes these remaining portions in subsequent iterations or batches [1]. To account for this, we ensure that our dataset reflects this practical constraint, making it more applicable not only for rigorous evaluations but also for practical scenarios: After generating 100 samples for each distribution, we ‘clip’ each sample to fall within the range of 1 to 500 for the bursttimes of each job. This process reflects resource allocation constraints inherent to CPUs. Values less than 1 are set to 1, representing the minimum resource allocation that a CPU can assign to a job. Similarly, values greater than 500 are set to 500, representing the maximum resource allocation that a CPU can assign to a job. This is how, from a set of randomly sampled distributions, we finalized these job clusters, adhering to the realistic constraints of CPU processing.
  2. 2.
    Normalisation of the Clusters: After the clusters are ready, we normalise all finalised samples. We divide each bursttime by the sum of the bursts present in the sample. After that we multiply each burst by 25000. We normalised each sample such that the sum of bursts in each sample across clusters are (almost) equal to 25000. By scaling the data, we eliminate the influence of magnitude disparities, allowing us to focus solely on the algorithmic performance with respect to bursttime distribution patterns between any pair of samples.
Furthermore, for a fine-grained comparison, we’ve used box plots for each metric to analysis across various distributions for each algorithm. We’ve referred to our algorithm as “FairBatch” here and have taken the ceil of mean of the bursttimes as the timeQuantum in RR as it remains almost equal for all the lists under examination.In diagram. We describe the legends used in experiments in table 1.
Abbreviation Description Used in
Nor Cluster Collection of arrays of jobs following the underlying Normal distribution Fig:2, 3, 4, 5
Expo Cluster Collection of arrays of jobs following the underlying Exponential distribution Fig:2, 3, 4, 5
Geo Cluster Collection of arrays of jobs following the underlying Geometric distribution Fig:2, 3, 4, 5
N.Bio Cluster Collection of arrays of jobs following the underlying Negative Binomial distribution Fig:2, 3, 4, 5
Psn Cluster Collection of arrays of jobs following the underlying Poisson distribution Fig:2, 3, 4, 5
Ufm Cluster Collection of arrays of jobs following the underlying Uniform distribution Fig:2, 3, 4, 5
Bimodal Cluster Collection of arrays of jobs following the underlying Bimodal distribution Fig:2, 3, 4, 5
Trimodal Cluster Collection of arrays of jobs following the underlying Trimodal distribution Fig:2, 3, 4, 5
Multimodal Cluster Collection of arrays of jobs following the underlying Multimodal distribution Fig:2, 3, 4, 5
Gamma Cluster Collection of arrays of jobs following the underlying Gamma distribution Fig:2, 3, 4, 5
Cauchy Cluster Collection of arrays of jobs following the underlying Cauchy (Lorentz) distribution Fig:2, 3, 4, 5
Pareto Cluster Collection of arrays of jobs following the underlying Pareto distribution Fig:2, 3, 4, 5
1 FCFS Fig:2, 3, 4, 5
2 FairBatch Fig:2, 3, 4, 5
3 SRTF Fig:2, 3, 4, 5
4 LRTF Fig:2, 3, 4, 5
5 RR Fig:2, 3, 4, 5
6 CFS (SCHED_BATCH) Fig:2, 3, 4, 5
Table 1: Abbreviations and descriptions of clusters and algorithms
4.2 Discussion
Each algorithm is tested on 100 test cases present in each cluster, we represent their aggregated results in the form of boxplots. Before analysis, we are presenting a concise introduction of boxplots for readers.
Box plots (Fig:6) are data visualization techniques that provide a concise summary of the distribution of a dataset. They consist of a box, which represents the interquartile range (IQR) containing the middle 50% of the data, with a line inside marking the median. Whiskers extend from the box to the minimum and maximum values within a certain range, often 1.5 times the IQR. Outliers beyond the whiskers are plotted individually.
Figure 6: Boxplot
  • [leftmargin=*]
  • Median (Q2): Middle value of the dataset.
  • Upper Quartile (Q3): Median of the upper half of the dataset.
  • Lower Quartile (Q1): Median of the lower half of the dataset.
  • Interquartile Range (IQR): Range containing the middle 50% of the data.
  • Upper Whisker: Maximum value within 1.5 times the IQR from Q3.
  • Lower Whisker: Minimum value within 1.5 times the IQR from Q1.
  • Outliers: Individual data points beyond the whiskers.
Interpreting Boxplots:
In performance analysis, the common practice is to focus on the median, IQR, and range [41, 42]:
  1. 1.
    Median: Less sensitive to outliers compared to the mean, providing a more reliable measure of central tendency.
  2. 2.
    IQR: A robust measure of the spread of the middle 50% of the data, less influenced by extreme values.
  3. 3.
    Range: Gives an indication of the overall spread of values in the dataset, despite being affected by outliers.
  4. 4.
    Ignoring Outliers: While outliers can offer insights into extreme data points, they can also distort the analysis. By focusing on the IQR, or median, we can mitigate the impact of outliers and obtain a more accurate understanding of the typical behavior of the data [41].
  • First Phase of the Analysis: In this phase, we investigate the algorithms against ‘Nor cluster’, ‘Expo cluster’, ‘Geo cluster’, ‘N.Bio cluster’, ‘Psn cluster’, ‘Ufm cluster’, ‘Gamma cluster’, ‘Cauchy cluster’, ‘Pareto cluster’. These comparisons are primarily made to simulate homogeneous yet diverse patterns of CPU loads. For our average turnaround and waiting time, the benchmarking algorithm is SRTF as mentioned. We compare the algorithms based on their comparative performance with respect to SRTF as demonstrated across 2,3,4. We ignore the outliers and relatively compare the performances based on the IQR, median, and range.
    If we discretely compare the average turnaround, waiting and response time for FairBatch (2 in the figures):
    1. 1.
      Out of 9 cases, FairBatch has achieved the least average turnaroundtime in 7 cases after SRTF. In the case of ‘Nor Cluster’, FairBatch has achieved the third least average turnaroundtime after FCFS (1 in figures) and SRTF. In the case of ‘Psn Cluster’, it comes after SRTF, FCFS, and SRTF respectively in 3,
    2. 2.
      Out of 9 cases, FairBatch has achieved the least average waitingtime in 7 cases after SRTF. In the case of ‘Nor Cluster’ in 3, FairBatch has achieved the third least average waitingtime after FCFS (1 in figures) and SRTF. In the case of ‘Psn Cluster’, it comes after SRTF, FCFS, and SRTF respectively in 3.
    3. 3.
      Out of 9 cases, FairBatch has achieved the least average responsetime in 7 cases. In the case of ‘Nor Cluster’ and ‘Psn Cluster’ in 3, FairBatch has achieved the second least average responsetime after LRTF (4 in figures). The only algorithm that ever exceeds FairBatch in responsetime is LRTF only in these 2 cases.
    Investigating the performance in the case of ‘Nor Cluster’
    Here, the underlying distribution is the normal distribution, where approximately 68% of the data falls within one standard deviation of the mean, about 95% falls within two standard deviations, and around 99.7% falls within three standard deviations of the mean [39]. So, in this cluster, the majority of the jobs’ burst times are centred around a single value (the mean). In other words, most jobs are very comparable in size (bursttimes). Now, FCFS, being a non-preemptive algorithm, completes jobs without any context switch. As a result, here in terms of turnaround and waiting time, FCFS excels. However, due to FCFS’s non-preemptive nature, it cannot respond to other jobs waiting in the queue before one job is completely executed, leading to the issue of the convoy effect [1]. FairBatch, in this case, preemptively switches contexts based on the fairness ratio, which prioritizes waitingtime for jobs as well as progress in currently under-executing jobs. This is evident when comparing the average responsetime of FairBatch and FCFS. FairBatch not only competitively prioritizes efficiency but also gives weightage to fairness in job selection.
    On the other hand, LRTF achieves the best responsetime in this cluster. LRTF, being a preemptive algorithm, at every instance chooses the longest available job, performs bare minimum execution, and moves on to the next largest job. As most of the jobs are similar in size, this procedure results in excessive preemption, providing a “response” to each job in the least time. While FCFS was completely focused on efficiently executing all the jobs one after another, causing severe convey in the system, LRTF aggressively responded to almost every available job with marginal progress, which drastically reduced its efficiency 3. FairBatch, on the other hand, strikes an excellent balance between efficiency and fairness that is unparalleled even in SRTF.
    Observing the trade-off Arose in the case of ‘Psn cluster’
    The Poisson distribution is primarily used to model the likelihood of an event occurring a certain number of times within a specific period. We have selected this discrete distribution to examine the occurrence of repetitive loads in CPUs [43]. Since batch processing is a CPU-bound process, repetitive loads are frequently observed in the system. Consequently, this cluster is heavily influenced by frequently occurring comparable jobs. FCFS, as a non-preemptive algorithm, aggressively executes repetitive jobs, causing other jobs to wait disproportionately and eventually starve. The next algorithm that outperforms FairBatch is CFS. As CFS (SCHED_BATCH) is a widely used finely tuned scheduling policy for large batches, operating on the principles of virtual runtime and nice values, it is well-suited for managing commonly observed repetitive loads on systems. However, due to its design [^4] https://elixir.bootlin.com/linux/v5.19.9/source/kernel/sched/fair.c##L7429 lacking a suitable preemption-supportive mechanism and heuristics, it does not perform well in terms of response time.
    On the other hand, LRTF aggressively preempts and responds to frequently repetitive loads. Similar to the ‘Nor Cluster’, while it achieves excellent responsiveness, it shows marginal efficiency. In contrast, FairBatch does not monopolize resources for repetitive jobs but evenly distributes them among all jobs while maintaining competitive efficiency. This demonstrates that the FairBatch algorithm exhibits strong adaptability and resilience in handling the randomness present in different data distributions.
  • Second Phase of Analysis: In the previous phase, we have observed the unparalleled superiority in terms of both efficiency and fairness of FairBatch across 7 out of 9 cases. We have also analysed when there is a harsh trade-off between efficiency and fairness, how all algorithm falls apart in consistently maintaining the trade-off except FairBatch. In order to delve into more on these scenarios comprising of critical trade-offs, after the first phase, here we primarily investigate bimodal, trimodal and multimodal clusters. If we discretely compare the average turnaround, waiting and response time for FairBatch in 5:
    1. 1.
      FairBatch achieves the least average turnaroundtime after FCFS and SRTF in all three cases.
    2. 2.
      FairBatch achieves the least average waitingtime after FCFS and SRTF in all three cases.
    3. 3.
      FairBatch achieves the least average responsetime in all three cases.
    In regular batch processing, heterogeneous loads are not commonly observed unless the batch is sufficiently large. These three distributions were created by concatenating more than one heterogeneous subpopulation of normal distributions. As a result, the central tendency of the measurements is inherently more complex than the previous clusters. These clusters are generated to adversarially test the algorithms to observe how they perform when there are multiple distinct subpopulations creating heterogeneous loads on the CPU. Depending on the sequence of the subpopulations, the arrival sequence of the jobs changes, potentially resulting in a change in the average turnaround, waiting, and response times.
    FCFS has a distinct advantage in these cases: it minimizes the maximum turnaroundtime across jobs for any finite arrival sequence of jobs [44]. Irrespective of the sequence of these subpopulations, FCFS achieves the least turnaround and waiting times across all three clusters after SRTF in our experiment. While it achieves superior efficiency due to its theoretical guarantees, it is evident from Figure 5 that it does not perform well in terms of responsetime in any of these clusters. It suffers from severe starvation and lags in fairness in job selection. FairBatch in all these cases performs excellently in handling the efficiency and fairness trade-off as evident from the results in figure 5. Not only FCFS, but also CFS and LRTF along with others do not perform well when it comes to selecting jobs in a fairer way while maintaining competitive efficiency unlike FairBatch.
  • Stability of the performance: another advantage of FairBatch
    FairBatch offers another distinct advantage over algorithms like Round Robin and FCFS due to its stability and predictability. While the sequence of execution and overall results of these algorithms heavily depend on the sequence in which the job arrived, FairBatch eliminates this variability by arranging and executing jobs based on fairnessratio and timeQuantum. This characteristic makes the algorithm a more reliable and consistent scheduling algorithm, ensuring stability and predictability in its performance.
After a comprehensive comparison of all algorithms across various distributions as demonstrated in fig:2, 3, 4, 5, it is evident that the FairBatch performs exceptionally competitively with respect to the benchmark (SRTF) for all distributions showcasing its superiority through the delicate balance it strikes between drastically reducing waitingtime, ensuring responsiveness, and eliminating starvation. With stability, predictability, and unmatched efficiency, our algorithm bridges the gap between suitable responsiveness and required efficiency through a fair and holistic approach, as demonstrated extensively.